OpenStudy (anonymous):

Area of a parallelogram a= [1,-4], = [3,-5]

OpenStudy (anonymous):

is there supposed to be a b = [3,-5]

OpenStudy (anonymous):

yup, typo *

OpenStudy (amistre64):

magnitude of the cross product i think is a way to determine it

OpenStudy (amistre64):

|a||b| sin(t); where t is the angle between a and b

OpenStudy (amistre64):

\[sin(t)=\frac{|axb|}{|a||b|}\] so yeah, the magnitude of the cross product

OpenStudy (anonymous):

@amistre64 how can i do the cross product when only x and y are given ?

OpenStudy (amistre64):

plug in z=0 :)

OpenStudy (anonymous):

ohhh! .. facepalm

OpenStudy (amistre64):

a= [1,-4], = [3,-5] x 1 3 y -4 -5 z 0 0 x = 0-0 -y = 0-0 z = -5+12 the magnitude of [0, 0, 7 ] = 7

OpenStudy (amistre64):

of course the cross of R^2 simplifies as the determinant

OpenStudy (anonymous):

Okay thanks ! :)

OpenStudy (anonymous):

wouldnt axb = 0 ? :S

OpenStudy (anonymous):

oh nvm

OpenStudy (amistre64):

my teacher did an R^2 cross a few semesters ago, and I was like, aint it just the determinant? and she was like or just plug in z=0 :)

OpenStudy (anonymous):

yeah i think its easier to plug in the zero :P

OpenStudy (amistre64):

have fun ;)